A bag contains 3 black balls and 2 red balls,a ball is picked at random and replaced.

what is the probability they are both black?


1
Expert's answer
2022-02-21T18:35:33-0500

Solution:

3 black balls and 2 red balls.

Total balls = 5

2 balls are picked with replacement.

The probability they are both black"=\dfrac{3}{5}\times \dfrac{3}{5}=\dfrac{9}{25}"